OpenStudy (ms_sp):

1. A rectangular flower box holds 1232in^3 of soil. Find the length of the box if it is 22 in. long and 7 in wide. 2. A flag is in the shape of a triangle and has an area of 6ft^2. Find the length of base if its height is 4ft.

OpenStudy (ms_sp):

@LannyXX @GIitter

OpenStudy (lannyxx):

Hmmm. I think it would be more effective for you to attempt those problems yourself, and if you are STUCK or DONT GET something, you explain what you did here and we can help! that way this would be more active on your part. part of getting better at problem solving is to experiment and investigate for yourself~

OpenStudy (ms_sp):

2L+2W= 2(w+7)=1232 correct?

OpenStudy (ms_sp):

@Will.H

OpenStudy (will.h):

Rectangular box volume is V = l * w * h l is length 2 is width and h i height and v is the volume now we know V = 1232 l = unknown h = 22 w = 7 1232 = l * 7 * 22 you can take it from here for the 1st one?

OpenStudy (ms_sp):

@Will.H : thank you for correcting what I did wrong, on number 1

OpenStudy (will.h):

as for the 2nd one triangle are formula is A = 1/2 * b * h A = 6 b = unknown h = 4 6 = 1/2 * b * 4 again you can take it from here? and you welcome

OpenStudy (will.h):

let me known if you need additional clarification

OpenStudy (ms_sp):

Yes I can, thank you for helping. Sorry, if I was bothering you. I will let you know.

OpenStudy (giitter):

Looks like you're covered. Im sorry I couldn't answer I was a little busy